OpenStudy (anonymous):

expand the log expression; logb(57/74)

OpenStudy (anonymous):

\[\log_b(\frac{A}{B})=\log_b(A)-\log_b(B)\]

OpenStudy (anonymous):

would that be under a radical ?

OpenStudy (jhannybean):

following @satellite73 's formula there, what is A and what is B in your problem?

OpenStudy (jhannybean):

And I don't believe it would be under a radical.

OpenStudy (anonymous):

errr the square root sign

OpenStudy (jhannybean):

I don't see a square root sign? O_o

OpenStudy (anonymous):

1/2logb57+1/2logb74 1/2logb57-1/2logb74 \[\sqrt{logb-\log74}\] logb1/2(57-74) are my answer choices

OpenStudy (jhannybean):

oh, is your problem \(\large \log_b \sqrt{\frac{57}{74}} \)

OpenStudy (anonymous):

yes

OpenStudy (jhannybean):

Oh okay. This can be solved like.. \[\large \log_b \left(\frac{A}{B}\right)^c = c \log_b A - c \log_b B\]

OpenStudy (jhannybean):

in this case, your "c" or your power will be 1/2 as we change the square root into a "power" format. Then we can just simply use our log identities.

OpenStudy (anonymous):

so 1/2logb57-1/2logb74

OpenStudy (jhannybean):

Yess good job :D

OpenStudy (anonymous):

but couldnt it alos be d?

OpenStudy (jhannybean):

mmhmm.

OpenStudy (jhannybean):

Nevermind. the "1/2" would come before the log b... the format of that solution statement is wrong. Therefore it cannot be D.

OpenStudy (anonymous):

oh yea i get it. ive never seen the 1/2 behind the log

OpenStudy (anonymous):

can you help me with one more thing?

OpenStudy (jhannybean):

\[\large \log_b\left(\frac{A}{B}\right)^c = c \cdot [\log_b (A) - \log_b (B)]\]

OpenStudy (jhannybean):

That means \(c\log(A) - c\log(B)\)

OpenStudy (jhannybean):

Sure thing.

OpenStudy (anonymous):

i know the answer is b or d but idk how to zoom in on my calculator to try to see the numbers

OpenStudy (jhannybean):

Which problem are we looking at? Unfortunately i cannot help you solve all of them because that would be violating the code of conduct of OS.

OpenStudy (anonymous):

oh im sorry i only meant to send you a screenshot but question 13

OpenStudy (jhannybean):

Alriiighty :)

OpenStudy (anonymous):

yea . idk how to zoom in

OpenStudy (anonymous):

here it is

OpenStudy (jhannybean):

Oh, you don't need to zoom in,you need to zoom out.

OpenStudy (anonymous):

but i have to see which equation so i have to see the #'s

OpenStudy (jhannybean):

Do you recall that \(\large \ln(x) = \frac{1}{x}\)?

OpenStudy (anonymous):

yes sir

OpenStudy (anonymous):

the answer is d?

OpenStudy (jhannybean):

So we can immediately eliminate the first choice,since that repr

OpenStudy (jhannybean):

represnts a parabola* and yes :) you are right.

OpenStudy (jhannybean):

http://fooplot.com/plot/5izc4llikx

OpenStudy (anonymous):

thank you !

OpenStudy (jhannybean):

no problemooo